OpenStudy (anonymous):

A sporting goods store sells right-handed and left-handed baseball gloves. In one month, 15 gloves were sold for a total revenue of $702. Right-handed gloves cost $42 and left-handed gloves cost $51. How many left-handed gloves did they sell?

OpenStudy (nikato):

Let l= left hand and r=right hand gloves

OpenStudy (nikato):

set up these two equations l+r=15 42r+51l= 702

OpenStudy (anonymous):

okay

OpenStudy (nikato):

and subtracting l on both sides on the first equation gives you r=15-l now substitute this into the second equation

OpenStudy (anonymous):

so itll be 42(15-l) +51l=702

OpenStudy (nikato):

correct. and now you want to solve for l. you can start with distributing the 42

OpenStudy (anonymous):

okay one moment please

OpenStudy (anonymous):

the left hand is 8?

OpenStudy (nikato):

corect

OpenStudy (anonymous):

and the right hand?

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

is it 7?

OpenStudy (nikato):

the question doesnt ask for right hand but if you want, to fidn the right, you just substitute 8 for l in one of the original equations to solve for r. and yes, 7 is corect